The gross domestic product, G, of Switzerland was 280 billion dollars in 2007. Give a formula for G (in billions of dollars) t years after 2007 if G increases by
3% a year

300 billion a year

Answer by Boreal(15235)

G=280(1.03)^n, where n is years after 2007.
300=280(1.03)^n
divide both sides by 280
1.0714=1.03^n
ln(1.0714)=n*ln (1.03)
n=ln (1.0714)/ln (1.03)=2.334 years, if this is the question desired.